Q6: Subnet the Class C IP Address 195.1.1.0 So that you have 10 subnets each with a maximum 12 hosts on each subnet. List the Address on host 1 on subnet 0,1,2,3,10.

Q7: Subnet the Class C IP Address 205.11.2.0 so that you have 30 subnets.

· What is the subnet mask for the maximum number of hosts?

· How many hosts can each subnet have?

· What is the IP address of host 3 on subnet 2?

Q8: You have been allocated a class A network address of 29.0.0.0. You need to create at least 20 networks and each network will support a maximum of 160 hosts. Would the following two subnet masks Work?

255.255.0.0 and or 255.255.255.0

Q9: Given the subnet Mask 255.255.255.192 What is the host address and subnet of the following

IP address 197.1.2.67.
Answer #1

Q6 answer:

Current mask= 255.255.255.0

Bits needs for 10 subnets =4 =24 =16 possible subnets

Bits needs for 12 hosts = 4 = 24 = 16-2=14 possible hosts.

So our mask in binary =11110000= 240 decimal

Final Mask =255.255.255.240

Q7 answer:

Current mask= 255.255.255.0

Bits needs for 30 subnets =5 =25 =32 possible subnets

Bits left for hosts = 3 = 23 = 8-2=6 possible hosts.

So our mask in binary =11111000= 248 decimal

Final Mask =255.255.255.248

Address of host 3 on subnet 2 is

subnet 2 =00010000 host 3 =000000011

Add the two together =00010011=19

therefore IP address of host 3 on subnet 2 =205.11.2.19

Q8 answer:

Yes both would work.

Mask 255.255.0.0 has 8 bits for the subnet and 16 bits for the host

8 bits would accommodate 28=256 subnets

16 bits would accommodate 216= over 64000 hosts

Mask 255.255.255.0 has 16 bits for the subnet and 8 bits of the host.

Have possible 28 -2 hosts =254 which is enough

Q9 answer:

192 in binary =11000000 gives 4 possible subnets of (showing 2 most significant bits):

00,01,10,11

67 in binary =01000011

So Applying Mask: